OpenStudy (anonymous):

solve 3tan2x-9=0 in interval [0,2pi)

OpenStudy (anonymous):

tan 2x = 9/3 = 3 = tan 71.57 deg. 2 answers within period (0 < x < 2Pi) x1 = 71.57/2 = 35.78, and x2 = 35.78 + 180 = 215.78 deg